# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

### v2

@@ IV Methods
All simulations are done in the NVT ensemble at room temperature $T=$298\text{\,}\mathrm{K}$\$ with the MB-Pol force field [94, 95, 96]. For starting configurations, 94 equally spaced initial configurations are extracted from a $18.8\text{\,}\mathrm{n}\mathrm{s}$ long simulation using the SPC/E force field [97] and the GROMACS molecular dynamics software [98]. The SPC/E simulations are done with a time step of $2\text{\,}\mathrm{f}\mathrm{s}$ and a velocity rescaling thermostat [99], with a relaxation time of $1\text{\,}\mathrm{p}\mathrm{s}$. Each of the initial configurations is equilibrated with the MB-Pol force field for $20\text{\,}\mathrm{p}\mathrm{s}$, and afterward, production runs with an average runtime of $0.86\text{\,}\mathrm{n}\mathrm{s}$ are executed using the LAMMPS software [100]. Again, a velocity-rescaling thermostat is employed. However, for MB-Pol simulations, the relaxation time is set to $5\text{\,}\mathrm{p}\mathrm{s}$, and the time step is $0.2\text{\,}\mathrm{f}\mathrm{s}$. The box size is $2\text{\,}\mathrm{n}\mathrm{m}$ in the x and y dimensions and $6\text{\,}\mathrm{n}\mathrm{m}$ in the z dimension. This size is chosen to ensure that the box contains more than twice as much air as water, since the water slab spans approximately $2.6\text{\,}\mathrm{n}\mathrm{m}$. The box is filled with $352$ water molecules and the two interfaces lie in the $xy$ plane. The total dipole moment of the simulation box is computed with the modified TTMF-4 model, which describes the electrostatics of the MB-Pol potential [64, 94]. The same model is used to predict the linear absorption profile, where we compute polarization density trajectories from monopole and dipole density trajectories. Molecular polarizabilities are computed with single-molecule quantum chemistry calculations using the software Gaussian 16 [101]. The theoretical level is CCSD(T)/aug-cc-pVTZ and B3LYP/aug-cc-pVTZ in the parameterization of $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$, appearing in Equations (19) and (20), respectively. Electrostatic interactions between the induced multipoles are computed as described in SM Sections V B and VIII [60], involving the Ewald summation algorithm included in OpenMM [102]. The simulation of bulk water is done in a cubic box with a length of $2.0727\text{\,}\mathrm{n}\mathrm{m}$ consisting of 297 molecules. The magnetic dipole contribution is extracted from a different trajectory, because it requires a much smaller write-out frequency. The initial configurations for the trajectories used to extract the bulk dielectric constant, molecular hyperpolarizability, and magnetic dipole contribution are taken from a trajectory generated with the SPC/E model after a simulation time of $20\text{\,}\mathrm{n}\mathrm{s}$ and $40\text{\,}\mathrm{n}\mathrm{s}$, respectively. The first $10\text{\,}\mathrm{p}\mathrm{s}$ of the MB-Pol trajectory is discarded, and a trajectory of length of $400\text{\,}\mathrm{p}\mathrm{s}$ is used to calculate the bulk dielectric constant and the molecular hyperpolarizability, and a trajectory of length of $300\text{\,}\mathrm{p}\mathrm{s}$ to calculate the second-order magnetic dipole susceptibility. The simulation parameters of the bulk simulations are the same as for the interface. We use the recommended settings of the MB-Pol potential from the GitHub repository "MBX - Version 0.7" from the Paesani group [103]. The write-out frequency is $3.2\text{\,}\mathrm{f}\mathrm{s}$ for all spectra except $0.4\text{\,}\mathrm{f}\mathrm{s}$ for the magnetic dipole contribution. In all simulations, periodic boundary conditions are imposed. However, we correct for the external field modification arising from the periodic replicas as described in SM Section IX [60], previously applied for the calculation of dielectric profiles [79, 68, 85]. The smoothing procedure is described in SM Section X [60].
V Acknowledgements
Appendix A Prediction of Locally Resolved SFG Spectra
@@
oscillating with the frequency $\omega^{\alpha}$, $\alpha\in\{\mathrm{SFG},\mathrm{VIS},\mathrm{IR}\}$ and $\delta_{ij}$ is the Kronecker delta. Hence, the symbol $\mathcal{F}^{\alpha}_{i}$ is a placeholder for the spatially constant amplitude of D-fields or E-fields on the relevant length scale, depending on the polarization. The external fields are given by $F^{\mathrm{\alpha}}_{i}(t)=\mathcal{F}^{\mathrm{\alpha}}_{i}e^{-i\omega^{\mathrm{\alpha}}t}+c.c.$}. By formulating the theory with respect to these fields, we avoid locality approximations. We note that $\tilde{s}^{(2)}_{ijk}(z)$ is distinct from the second-order susceptibility $\tilde{\chi}^{(2)}_{ijk}(z)$, which defines the response to electric E-fields. Here, we only consider the signal arising from the interface on the length scale at which the external fields are spatially constant. External field gradients lead to so-called bulk quadrupole contributions [41, 25, 42], which can be experimentally estimated using combined transmission-reflection SFG [105] or combined SFG/DFG techniques [21, 50] and are discussed in SM Section IV C [60]. The SFG spectrum follows by integration over $z$ as
@@
where $\Delta k_{z}$ is the wave vector mismatch defined in SM Section II [60]. We emphasize that Equation (11) holds whenever the external fields can be assumed to be constant in the region where the dielectric profile is inhomogeneous. This assumption is well satisfied in our case, because the shortest wavelength considered is $\lambda^{\mathrm{SFG}}\approx$600\text{\,}\mathrm{n}\mathrm{m}$\$}, and the region where the dielectric profile is inhomogeneous is only a few angstroms thick, as demonstrated in this work. Here we utilize the additional approximation $e^{-i\Delta k_{z}z}\tilde{s}^{(2)}_{ijk}(z)\approx\tilde{s}^{(2)}_{ijk}(z)$. At the air waiter interface, this is a good approximation, as a typical value for $\Delta k_{z}$ is $0.02\text{\,}\mathrm{n}\mathrm{m}^{-1}$, whereas the thickness of the air-water interface is about $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ [21]. A derivation of Equations (7)-(11) is provided in SM Sections I and II [60].
Appendix B Molecular Multipole Contributions
@@
While the total current $j^{(2)}_{i}(z,t)$ does not depend on the choice of the molecular origin, the individual contributions in Equation (12) do [108, 74]. To determine the optimal position of the molecular origin, we consider the SFG signal from an interface with isotropically oriented molecules $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}$, which we call an isotropic interface. This isotropic interface is created by cutting bulk water at an arbitrary $z$ position into two halves. Consequently, boundary contributions created by the change in density at the interface are present, but the molecules’ orientational distribution is inversion symmetric. A similar interface was created to predict the Bethe potential [109]. We test three molecular origins for the calculation of multipoles and find that if we choose the molecular center of mass as the molecular origin, $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}\approx\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$} does hold in good approximation. Thus, choosing the molecular center of mass as the expansion center ensures that the electric dipole contribution is solely due to molecular orientation, as discussed in detail in SM Section VI [60]. In contrast, choosing the molecular center different from the center of mass introduces significant boundary contributions in the electric dipole contribution $\tilde{S}^{(2,\mathrm{D})}_{ijk}$. Moreover, the center of mass maximizes the decoupling of molecular translations, vibrations, and rotations [110, 111, 112]. The theory of our multipole decomposition is described in SM Sections III and IV [60]. There, we also derive the constitutive relations, including nonlinear multipolar source terms, using the Lorentz field approximation in planar geometry, based on works by Mizrahi and Sipe [113] and Hirano and Morita [42]. These equations are solely needed to aid the interpretation of experimental spectra.
Appendix C Fluctuation-Dissipation Relations within the Off-Resonant Approximation
We assume that the VIS field interacts off-resonantly with the system, meaning it does polarize the molecules but does neither excite higher electronic levels nor influence the nuclei trajectories. This approximation is valid since the VIS field oscillates too rapidly to influence the motion of the nuclei yet slowly enough to act adiabatically on the distribution of the electrons. In this limit the SFG signal arises from a first-order perturbation expansion with respect to the external field from the IR light source [114]. The molecular multipoles are determined by the equations
@@ Appendix C Fluctuation-Dissipation Relations within the Off-Resonant Approximation
where $\mu^{n}_{i}(t)$ and $Q^{n}_{ij}(t)$ are the induced electric dipole and electric quadrupole moments of the nth-molecule, and $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$ are the electric dipole and the electric quadrupole polarizabilities of the $n^{\mathrm{th}}$ molecule. The local field factor $f_{ij}^{n}(t)$ transforms an external field $F_{i}(t)$ into the local E-field acting on the nth molecule, $E^{n}_{i}(t)=f_{ij}^{n}(t)F_{j}(t)$}, and is determined in a self-consistent manner such that the multipoles induced by $E^{n}_{i}(t)$ produce the field $E^{n}_{i}(t)-F_{i}(t)$. It is demonstrated in SM Section V B [60] that it is incorrect to solve separate self-consistent field equations for the VIS and the SFG field in the time domain, as proposed previously [114, 115, 116]. Rather, if a timescale-separation approximation is applied, Equations (19) and (20) each split into components at SFG and VIS frequencies, as described in SM Section V B [60]. The polarization contributions of the second-order response profiles, defined in Equation (17), are determined by a first-order perturbation expansion, for which we introduce the fluctuation-dissipation relations
@@
where $\langle$$\rangle$ denotes ensemble averaging, $T$ is the temperature, $k_{B}$ the Boltzmann constant, $\theta(x)$ denotes the Heaviside function and the z-position of the $n^{\mathrm{th}}$ molecule is denoted as $z^{n}$. These expressions omit the off-resonant hyperpolarizability, which can be included as shown in SM Section V [60], but does not contribute to the imaginary part of SFG spectra. Equations (21) and (22) follow from time-dependent perturbation theory using the perturbation Hamiltonian $H_{\mathrm{int}}(t)=-P_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P_{i}$ is the total system’s dipole moment. Hirano and Morita [74] suggest the perturbation expansion using the modified perturbation Hamiltonian $H_{\mathrm{int}}^{+}(t)=-P^{+}_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P^{+}_{i}(t)$ is the dipole moment of the upper half of the simulation box. It is demonstrated in SM Section V D [60] that this introduces a spurious contribution from the interface between the two halves. We note that while we treat electric dipole and electric quadrupole contributions in a precise manner, we approximate magnetic dipole contributions by considering only the leading order term of an expansion in terms of electric multipole moments, as outlined in SM Section V C [60]. We do not compute the position-resolved profile of magnetic dipole contributions, but only the overall contribution to the SFG signal, which can be extracted from a simulation of a bulk system, according to a fluctuation-dissipation relation presented in SM Section V C [60]. Hence, we omit magnetic dipole contributions, whenever we present the response profiles $\tilde{s}^{(2)}_{ijk}(z)$, but consider them when we present the total spectra $\tilde{S}^{(2)}_{ijk}$. In contrast to the electric dipole and electric quadrupole contributions, the magnetic dipole contribution does depend on $\omega^{\mathrm{VIS}}$}, set to $2730\text{\,}\mathrm{T}\mathrm{H}\mathrm{z}$, which is the center frequency of the VIS-field employed in the experimental measurement of the bending band [50] we compare with. The VIS frequency in the experimental reference SFG spectra of the stretching band [31, 117] is quite similar.
Appendix D Linear Dielectric and Absorption Profiles
@@
The extraction of $\tilde{s}^{(1,\mathrm{P})}_{ij}(z)$ from molecular dynamics simulation has been described before [80, 81, 82, 83, 84, 85, 86] and is reproduced in SM Sections V B [60]. At the interface, the dielectric profile is tensorial, and the component perpendicular to the interface is given by
@@
Further information about the dielectric profiles is given in SM Section XI [60].
Appendix E Electric Dipole SFG Contribution as a Fingerprint for Interfacial Structure

### v3

@@ IV.1 Technical Details
The simulation of bulk water is done in a cubic box with a length of $2.07\text{\,}\mathrm{n}\mathrm{m}$ consisting of 297 molecules. The magnetic dipole contribution is extracted from a different trajectory, because it requires a much smaller write-out frequency. The initial configurations for the trajectories used to extract the bulk dielectric constant, molecular hyperpolarizability, and magnetic dipole contribution are taken from a trajectory generated with the SPC/E model after a simulation time of $20\text{\,}\mathrm{n}\mathrm{s}$ and $40\text{\,}\mathrm{n}\mathrm{s}$, respectively. The first $10\text{\,}\mathrm{p}\mathrm{s}$ of the MB-Pol trajectory is discarded, and a trajectory of length of $400\text{\,}\mathrm{p}\mathrm{s}$ is used to calculate the bulk dielectric constant and the molecular hyperpolarizability, and a trajectory of length of $300\text{\,}\mathrm{p}\mathrm{s}$ to calculate the second-order magnetic dipole susceptibility. The simulation parameters of the bulk simulations are the same as for the interface. We use the recommended settings of the MB-Pol potential from the GitHub repository "MBX - Version 0.7" from the Paesani group [104]. The write-out frequency is $3.2\text{\,}\mathrm{f}\mathrm{s}$ for all spectra except $0.4\text{\,}\mathrm{f}\mathrm{s}$ for the magnetic dipole contribution. In all simulations, periodic boundary conditions are imposed. However, we correct for the external field modification arising from the periodic replicas as described in SI Section IX. The smoothing procedure and the error estimation method of the predicted spectra are explained in SI Section X. Details on the comparison between the experimental and simulated absolute spectra presented in Figure 1 (k) are given in SI Section XI. Further references are cited in the SI [15, 54, 97, 136, 129, 67, 122, 60, 91, 42, 18, 142, 118, 77, 28].
IV.2 Prediction of Locally Resolved SFG Spectra
@@
oscillating with the frequency $\omega^{\alpha}$, $\alpha\in\{\mathrm{SFG},\mathrm{VIS},\mathrm{IR}\}$ and $\delta_{ij}$ is the Kronecker delta. Hence, the symbol $\mathcal{F}^{\alpha}_{i}$ is a placeholder for the spatially constant amplitude of D-fields or E-fields on the relevant length scale, depending on the polarization. The external fields are given by $F^{\mathrm{\alpha}}_{i}(t)=\mathcal{F}^{\mathrm{\alpha}}_{i}e^{-i\omega^{\mathrm{\alpha}}t}+c.c.$}. By formulating the theory with respect to these fields, we avoid locality approximations. We note that $\tilde{s}^{(2)}_{ijk}(z)$ is distinct from the second-order susceptibility $\tilde{\chi}^{(2)}_{ijk}(z)$, which defines the response to electric E-fields. Several studies have defined $\tilde{\chi}^{(2)}_{ijk}(z)$ using constitutive relations that involve an position-independent effective interfacial dielectric constant $\tilde{\varepsilon}^{\alpha}_{\mathrm{eff}}$ [31, 140, 141]. Within that framework $\tilde{s}^{(2)}_{ijk}(z)$ and $\tilde{\chi}^{(2)}_{ijk}(z)$ are related by
$\tilde{\chi}^{(2)}_{yyz}(z)$ $=\tilde{\varepsilon}^{\mathrm{IR}}_{\mathrm{eff}}\tilde{s}^{(2)}_{yyz}(z)$ (12)
$\tilde{\chi}^{(2)}_{zzz}(z)$ $=\tilde{\varepsilon}_{\mathrm{eff}}^{\mathrm{SFG}}\tilde{\varepsilon}_{\mathrm{eff}}^{\mathrm{VIS}}\tilde{\varepsilon}_{\mathrm{eff}}^{\mathrm{IR}}\tilde{s}^{(2)}_{zzz}(z)\,$ (13)
with similar relations for other tensor elements. The relation in Equation (1) used by us is exact on the linear response level and has the advantage that it fully captures position-dependent dielectric effects.
In this work, we only consider the signal arising from the interface on the length scale at which the external fields are spatially constant, therefore we neglect contributions induced by external field gradients. These are so-called bulk quadrupole contributions [117, 88, 59], which can be experimentally estimated using combined transmission-reflection SFG [124], combined SFG/DFG techniques [39, 41] or configuration analysis and are discussed in SI Section IV C. It can be shown that $\tilde{S}^{(2)}_{yyz}$ is largely insensitive to bulk quadrupole contributions, provided that the transmitted-beam wavevectors remain approximately parallel [117]. Therefore, $\tilde{S}^{(2)}_{yyz}$ presented in Figure 1 (c), (g) and (k) can be assumed to be not significantly perturbed by bulk quadrupole contributions. However, this is not the case for $\tilde{S}^{(2)}_{zzz}$ presented in Figure 1 (d) & (h). We compare our theoretical prediction with a previous experimental configuration analysis [44] in SI Section XI and thereby confirm that our comparison is not significantly affected by bulk quadrupole contributions.
The SFG spectrum follows by integration of the second-order response profile over $z$ as
@@
where $\Delta k_{z}$ is the wave vector mismatch defined in SI Section II. We emphasize that Equation (14) holds whenever the external fields can be assumed to be constant in the region where the dielectric profile is inhomogeneous. This assumption is well satisfied in our case, because the shortest wavelength considered is $\lambda^{\mathrm{SFG}}\approx$600\text{\,}\mathrm{n}\mathrm{m}$\$}, and the region where the dielectric profile is inhomogeneous is only a few angstroms thick, as demonstrated in this work. Here we utilize the additional approximation $e^{-i\Delta k_{z}z}\tilde{s}^{(2)}_{ijk}(z)\approx\tilde{s}^{(2)}_{ijk}(z)$. At the air waiter interface, this is a good approximation, as a typical value for $\Delta k_{z}$ is $0.02\text{\,}\mathrm{n}\mathrm{m}^{-1}$, whereas the thickness of the air-water interface is about $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ [39]. A derivation of Equations (8)-(14) is provided in SI Sections I and II.
IV.3 Molecular Multipole Contributions
@@
The second-order response profile is decomposed in analogy to Equation (15) into
@@
While the total current $j^{(2)}_{i}(z,t)$ does not depend on the choice of the molecular origin, the individual contributions in Equation (15) do [27, 58]. To determine the optimal position of the molecular origin, we consider the SFG signal from an interface with isotropically oriented molecules $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}$, which we call an isotropic interface. This isotropic interface is created by cutting bulk water at an arbitrary $z$ position into two halves. Consequently, boundary contributions created by the change in density at the interface are present, but the molecules’ orientational distribution is inversion symmetric. A similar interface was created to predict the Bethe potential [101]. We test three molecular origins for the calculation of multipoles and find that if we choose the molecular center of mass as the molecular origin, $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}\approx\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$} does hold in good approximation. Thus, choosing the molecular center of mass as the expansion center ensures that the electric dipole contribution is solely due to molecular orientation, as discussed in detail in SI Section VI. In contrast, choosing the molecular center different from the center of mass introduces significant boundary contributions in the electric dipole contribution $\tilde{S}^{(2,\mathrm{D})}_{ijk}$. Moreover, the center of mass maximizes the decoupling of molecular translations, vibrations, and rotations [36, 57, 135].
The theory of our multipole decomposition is described in SI Sections III and IV. There, we also derive the constitutive relations, including nonlinear multipolar source terms, using the Lorentz field approximation in planar geometry, based on works by Mizrahi and Sipe [82] and Hirano and Morita [59]. These equations are solely needed to aid the interpretation of experimental spectra.
IV.4 Fluctuation-Dissipation Relations within the Off-Resonant Approximation
We assume that the VIS field interacts off-resonantly with the system, meaning it does polarize the molecules but does neither excite higher electronic levels nor influence the nuclei trajectories. This approximation is valid since the VIS field oscillates too rapidly to influence the motion of the nuclei yet slowly enough to act adiabatically on the distribution of the electrons. In this limit the SFG signal arises from a first-order perturbation expansion with respect to the external field from the IR light source [87].
The molecular multipoles are determined by the equations
@@ IV.4 Fluctuation-Dissipation Relations within the Off-Resonant Approximation
where $\mu^{n}_{i}(t)$ and $Q^{n}_{ij}(t)$ are the induced electric dipole and electric quadrupole moments of the nth-molecule, and $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$ are the electric dipole and the electric quadrupole polarizabilities of the $n^{\mathrm{th}}$ molecule. The local field factor $f_{ij}^{n}(t)$ transforms an external field $F_{i}(t)$ into the local E-field acting on the nth molecule, $E^{n}_{i}(t)=f_{ij}^{n}(t)F_{j}(t)$}, and is determined in a self-consistent manner such that the multipoles induced by $E^{n}_{i}(t)$ produce the field $E^{n}_{i}(t)-F_{i}(t)$. It is demonstrated in SI Section V B that it is incorrect to solve separate self-consistent field equations for the VIS and the SFG field in the time domain, as proposed previously [87, 63, 64]. Rather, if a timescale-separation approximation is applied, Equations (22) and (23) each split into components at SFG and VIS frequencies, as described in SI Section V B.
The polarization contributions of the second-order response profiles, defined in Equation (20), are determined by a first-order perturbation expansion, for which we introduce the fluctuation-dissipation relations
@@
where $\langle$$\rangle$ denotes classical ensemble averaging, $T$ is the temperature, $k_{B}$ is the Boltzmann constant, $\Theta(x)$ is the Heaviside function, $\delta(x)$ is the Dirac delta distribution, and $z^{n}$ is the z-position of the $n^{\mathrm{th}}$ molecule. These expressions omit the off-resonant hyperpolarizability, which can be included as shown in SI Section V, but does not contribute to the imaginary part of SFG spectra.
Equations (24) and (25) follow from time-dependent perturbation theory using the perturbation Hamiltonian $H_{\mathrm{int}}(t)=-P_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P_{i}$ is the total system’s dipole moment. Hirano and Morita [58] suggest the perturbation expansion using the modified perturbation Hamiltonian $H_{\mathrm{int}}^{+}(t)=-P^{+}_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P^{+}_{i}(t)$ is the dipole moment of the upper half of the simulation box. It is demonstrated in SI Section V D that this introduces a spurious contribution from the interface between the two halves.
It is shown in SI Section XII that Equations (24) and (25) provide a leading-order harmonic approximation of the quantum mechanical SFG response profile and that no quantum-correction factor needs to be applied. The harmonic quantum-correction factor $Q_{\mathrm{HA}}=\frac{\hbar\omega\beta}{1-e^{-\hbar\omega\beta}}$, which appears in the literature [98, 7, 128], relates quantum and classical time-correlation functions [98], but does not relate classical and quantum response functions.
We note that while we treat electric dipole and electric quadrupole contributions in a precise manner, we approximate magnetic dipole contributions by considering only the leading order term of an expansion in terms of electric multipole moments, as outlined in SI Section V C. We do not compute the position-resolved profile of magnetic dipole contributions, but only the overall contribution to the SFG signal, which can be extracted from a simulation of a bulk system, according to a fluctuation-dissipation relation presented in SI Section V C. Hence, we omit magnetic dipole contributions, whenever we present the response profiles $\tilde{s}^{(2)}_{ijk}(z)$, but consider them when we present the total spectra $\tilde{S}^{(2)}_{ijk}$. In contrast to the electric dipole and electric quadrupole contributions, the magnetic dipole contribution does depend on $\omega^{\mathrm{VIS}}$}, set to $2730\text{\,}\mathrm{T}\mathrm{H}\mathrm{z}$, which is the center frequency of the VIS-field employed in the experimental measurement of the bending band [41] we compare with. The VIS frequency in the experimental reference SFG spectra of the stretching band [31, 140] is quite similar.
IV.5 Linear Dielectric and Absorption Profiles
@@
The extraction of $\tilde{s}^{(1,\mathrm{P})}_{ij}(z)$ from molecular dynamics simulation has been described before [19, 20, 21, 47, 75, 76, 12] and is reproduced in SI Sections V B. At the interface, the dielectric profile is tensorial, and the component perpendicular to the interface is given by
@@
Further information about the dielectric profiles is given in SI Section XIII.
IV.6 Electric Dipole SFG Contribution as a Fingerprint for Interfacial Structure